### v2

@@ 5 Conclusion
  • A. Preliminaries
    • A.1. Overall Notations
    • A.2. MLLM Reasoning
    • A.3. Monte Carlo Tree Search
    • A.4. Verification Methods
  • B. Algorithm Details
    • B.1. Action Space
    • B.2. Reward Value in MCTS
  • C. More Details about Experimental Setup
    • C.1. Benchmarks and Datasets
    • C.2. Baselines
    • C.3. Implementation Details
  • D. Supplementary Results
    • D.1. Detailed Results on Multimodal Reasoning Benchmarks
    • D.2. Comparison with Strong Baselines
    • D.3. Integration with SFT
    • D.4. Weak-to-Strong Generalization
  • E. Case Study
Appendix A Preliminaries
This section describes overall notations (A.1), MLLM reasoning (A.2), monte carlo tree search (A.3), and verification methods (A.4).
A.1 Overall Notations
@@
A.2 MLLM Reasoning
With the advancement of computational resources and expanded datasets, MLLMs have demonstrated remarkable capabilities across various multimodal tasks Yin et al. (2023); Zhang et al. (2024c); Wang et al. (2024f). These range from visual reasoning tasks like chart understanding and visual question-answering Masry et al. (2022); Hartsock & Rasool (2024) to more complex perception-action tasks in autonomous driving and robotics Driess et al. (2023); Cui et al. (2024); Yang et al. (2025), where models must integrate visual inputs with decision-making processes. Central to these achievements is the development of effective reasoning methods, which can substantially enhance MLLM problem-solving capabilities, enabling even smaller models to achieve sophisticated reasoning abilities Wang et al. (2024b); Dong et al. (2024a); Thawakar et al. (2025). To formalize this reasoning process, we consider an autoregressive pre-trained MLLM serving as a policy model $\pi_{\theta}$. Given an input problem $x_{t}$, the model generates a sequence $output=(s_{0},s_{1},s_{2},...,s_{T})$ through iterative token prediction, where $s_{0}:=x_{t}$ represents the initial state and $s_{T}$ corresponds to the solution $y_{p}$. We define this generated sequence $(s_{0},s_{1},s_{2},...,s_{T})$ as a reasoning trajectory $y_{t}$. The conditional probability distribution of generating the complete reasoning trajectory is:
@@ A.2 MLLM Reasoning
Following Hao et al. (2023); Ge et al. (2024), we can conceptualize MLLMs as world models, with the complex reasoning process formulated as a Markov decision process. Specifically, when addressing complex reasoning challenges in real-world scenarios, at each time step $t$, the model receives a state $S_{t-1}$, comprising the original input problem $x$ and preceding reasoning steps $(s_{0},s_{1},s_{2},...,s_{t-1})$. The policy model $\pi_{\theta}$ then generates the current action $a_{t}=\pi_{\theta}(\Phi(S_{t-1}))$, which prompts the MLLM to produce the next reasoning step $s_{t}$. The entire process, from the initial step $s_{0}$ to the final output $s_{T}$, naturally forms a complete trajectory or chain of thought.
@@
Table 6: Method Comparison: Recent Multimodal Structured Reasoning Approaches.
Method Open-Source Only Data Volume Training-Free Training Cost $\downarrow$ Inference Cost $\downarrow$
AR-MCTS Dong et al. (2024a) 34.5K - High
Mulberry Yao et al. (2024) 260K High Low
LLaVA-CoT Xu et al. (2024) 100K High Low
URSA Luo et al. (2025) 1100K High Low
LlamaV-o1 Thawakar et al. (2025) 118K High Low
AStar (Ours) 0.5K Low Low
@@
Early work explored multimodal chain-of-thought (CoT) Zhang et al. (2024g), addressing the limitations of conventional MLLMs that typically defaulted to simple ’direct prediction’ mode due to the scarcity of high-quality long-chain reasoning data. However, these approaches often exhibit instability in multimodal reasoning tasks due to the distribution shift between training and inference Wang et al. (2024e). Two primary methods have emerged to address this challenge: explicit search mechanisms and teacher-guided distillation. Recent advances in these approaches are summarized in Table 6.
@@
(1) Explicit search mechanisms.   These approaches leverage explicit search structures (e.g., Monte Carlo tree search, MCTS) with specialized reward models to guide the exploration of solution paths. AR-MCTS Dong et al. (2024a) combines MCTS with active retrieval to expand the solution space and enhance performance, though its extensive iterations demand significant computational resources. Moreover, the quality and relevance of retrieved examples at each exploration step substantially impact model performance. Similarly, Mulberry Yao et al. (2024) employs collective knowledge to enable collaborative conjecture, search, and identification of effective reasoning paths via MCTS. However, it requires substantial training data (260K samples) generated with expensive proprietary models, making it computationally intensive.
(2) Teacher-guided distillation.   This approach focuses on distilling structured reasoning patterns through long-form CoT Wei et al. (2022); Zhang et al. (2024g); Luo et al. (2025) instruction data, typically requiring supervision from proprietary models for data synthesis. For instance, LLaVA-CoT Xu et al. (2024) explicitly structures visual reasoning into four stages (Summary, Caption, Reasoning, and Conclusion), utilizing 100K synthetic data samples for downstream task fine-tuning. Similarly, URSA Luo et al. (2025) proposes a three-module data synthesis strategy integrating CoT distillation, trajectory rewriting, and format standardization, which relies on Gemini-Flash-002.
A.3 Monte Carlo Tree Search
As a heuristic search algorithm, MCTS has demonstrated remarkable success in complex reasoning and decision-making environments Chaslot et al. (2008); Ye et al. (2021); Zhou et al. (2024). The algorithm conceptualizes search spaces as tree structures and has achieved significant breakthroughs across various domains, most notably in game-playing AI such as AlphaGo and AlphaZero Dong et al. (2020). The basic MCTS algorithm involves an iterative search process with four key steps: selection, expansion, simulation, and backpropagation. As an example in multimodal mathematical reasoning, Figure 10 illustrates the four phases in an iteration, expanding the tree and then updating reward values.
@@ A.3 Monte Carlo Tree Search
Figure 10: An illustration of the four phases in an iteration of MCTS for complex reasoning tasks (A.3).
@@
Leveraging MCTS, recent approaches like AR-MCTS Dong et al. (2024a) exploit MLLMs’ intrinsic capabilities and active retrieval for iterative exploration to enhance visual complex reasoning. However, these methods demand substantial computational resources due to extensive iterations. While approaches like Mulberry Yao et al. (2024) attempt to integrate tree structures through model training, they require large-scale datasets and significant computational overhead. Moreover, current methods typically employ static reasoning processes, lacking the ability to adapt reasoning strategies based on problem complexity. In contrast, our approach employs MCTS only during the generation of prior reasoning patterns (referred to as ”thought cards” in Sec. 3.2) and references these thought cards during inference to achieve efficient reasoning. This design enables AStar to adaptively match reasoning strategies to problem complexity, significantly reducing time complexity compared to traditional tree search methods while maintaining comprehensive search coverage and performance, thus achieving an optimal efficiency-effectiveness trade-off.
A.4 Verification Methods
@@
After obtaining multiple valid candidate solutions for the test multimodal query $x_{t}=[q;i]$, selecting the most accurate reasoning trajectory among candidates presents a critical challenge. In multimodal reasoning, self-consistency Wang et al. (2023) represents a simple yet effective approach, where the final answer is determined through majority voting among sampled paths:
@@
Recent advances in LLM reasoning verification have demonstrated significant improvements in reasoning capabilities through verification modules Wang et al. (2024d); Zhang et al. (2024b). This has motivated the integration of outcome and process supervision in multimodal domains Dong et al. (2024a); Luo et al. (2025) to enhance reasoning performance. However, the scarcity of multimodal supervision data and challenges in ensuring supervision signal quality often compromise verification effectiveness, with few high-performance open-source verification models currently available. Therefore, for simplicity, we leverage mature text-domain outcome reward models (ORM).
Specifically, in the text domain, given a text problem $q$ and its reasoning trajectory $y_{t}$, ORM ($\mathcal{D}\times\mathcal{S}\rightarrow\mathbb{R}$) assigns a single real-value to indicate whether $y_{t}$ is correct. ORM is usually trained with a cross-entropy loss Cobbe et al. (2021); Li et al. (2023):
@@
where $y_{g}$ is the golden answer ($y_{g}=1$ if $traj$ is correct else $y_{g}=0$), and $r_{traj}$ is the sigmoid score of $traj$ assigned by ORM. The effectiveness of ORM heavily depends on training data quality. For reasoning problems with definitive answers, researchers can construct the ORM training set automatically by: (1) sampling candidate solutions, and (2) labeling solutions based on answer correctness. While this approach may introduce false positives (incorrect reasoning leading to correct answers), prior work has demonstrated its effectiveness in training robust ORM models Lightman et al. (2024); Wang et al. (2024c). This enables the development of high-performing open-source verification models in the text domain.
To leverage these powerful text-domain verification models for visual reasoning while maintaining our principle of balancing performance and efficiency, we directly employ an MLLM to translate the original visual problems into pure text format.
@@
where $\text{prompt}_{translate}$ denotes a translation prompt (e.g., “please translate the question and visual image into a pure text format”). The resulting text question $q$ and the derived language-based reasoning steps derived from prior visual reasoning are then jointly processed by the ORM model for path selection and verification.
Table 7: Comparison with other search-based LLM and MLLM methods. Note that, most method contain limited space. In contrast, we define a rich set of reasoning actions, thus enhancing the upper bound of model reasoning capabilities.
@@
LLM Tree-of-Thought Yao et al. (2023) $a_{3}$: one-step thought
RAP Hao et al. (2023) $a_{5}$: divide and conquer
ReST-MCTS* Zhang et al. (2024b) $a_{3}$: one-step thought
MCTSr Zhang et al. (2024a) $a_{4}$: chain-of-thought, $a_{6}$: self-reflection
MLLM AR-MCTS Dong et al. (2024a) $a_{3}$: one-step thought
LLaVA-CoT Xu et al. (2024) $a_{4}$: chain-of-thought
URSA Luo et al. (2025) $a_{4}$: chain-of-thought
@@
Our experiments indicate that this approach achieves improvements over simple self-consistency with minimal computational overhead. Future work could explore integration with high-quality verification models for potentially better results. In this paper, we utilize off-the-shelf ORM model: Llama3.1-8B-ORM-Mistral-Data [^1] https://huggingface.co/RLHFlow/Llama3.1-8B-ORM-Mistral-Data.
Appendix B Algorithm Details
B.1 Action Space
Reasoning capabilities are fundamental for handling diverse tasks, encompassing multiple cognitive dimensions. Prior work has categorized reasoning into deductive, inductive, abductive, and analogical approaches Wang et al. (2024f). While open-source MLLMs demonstrate basic task competency, advanced models necessitate more sophisticated human-like reasoning abilities to achieve their full potential Zeng et al. (2024). A key insight from recent studies suggests that model reasoning performance is bounded by their available reasoning actions Wu et al. (2024b). Current frameworks, however, operate within a constrained action space (Table 7), which may limit the reasoning capabilities of MLLMs. To expand models’ reasoning capabilities beyond these constraints, we propose a comprehensive framework with six fundamental reasoning actions:
  • Visual Parsing ($a_{1}$): Extracting and analyzing visual information from input images to support multimodal reasoning tasks.
  • System Analysis ($a_{2}$): Analyzing the overall structure of the problem and identifying the constraints and conditions before addressing it, thereby clarifying task requirements effectively.
  • One-Step Thought ($a_{3}$): Generating the next one-step thought based on the given question and the preceding reasoning steps.

### v3

@@ 5 Conclusion
  • A. TL;DR: Main Contributions and Takeaways
  • B. Preliminaries
    • B.1. Overall Notations
    • B.2. MLLM Reasoning
    • B.3. Monte Carlo Tree Search
    • B.4. Verification Methods (Self-Consistency and Outcome Reward Model)
  • C. Algorithm Details
    • C.1. Action Space
    • C.2. Reward Value in MCTS
  • D. Detailed Experimental Setup
    • D.1. Benchmarks and Datasets
    • D.2. Comparison Baselines
    • D.3. Implementation Details
  • E. Supplementary Results
    • E.1. Detailed Results on Multimodal Reasoning Benchmarks
    • E.2. Comparison with Strong Baselines
    • E.3. Weak-to-Strong Generalization
    • E.4. Statistical Validation
  • F. Case Study
  • G. Further Discussions
Appendix A TL;DR: Main Contributions and Takeaways
Takeaways 1. External reasoning guidelines unlock latent capabilities in resource-efficient models, effectively bridging the performance gap with significantly larger architectures. (Figure 1,4). 2. The universality of high-level thought patterns transcends specific tasks, domains, and modalities, suggesting fundamental cognitive principles that generalize across diverse problem spaces (§3.2). 3. Minimal data requirements (500 samples) demonstrate the power of explicit pattern extraction over large-scale implicit learning approaches (§3.3). 4. AStar’s modular design enables seamless integration with post-training techniques like GRPO, creating complementary performance gains without requiring architectural modifications (Figure 7).
In this work, we focus on exploring how to enhance MLLMs’ complex visual reasoning capabilities that demand both visual perception and high-level cognition. While mainstream multimodal reasoning approaches typically rely on computationally intensive explicit search mechanisms or post-training procedures requiring substantial data and computational resources, no prior work has sufficiently explored combining external reasoning patterns with internal capabilities to improve both efficiency and effectiveness. To bridge this gap, we present AStar, a novel thought-augmented reasoning framework with the following contributions:
  • Novel Reasoning Paradigm: We introduce “thought cards”, a lightweight library of high-level reasoning patterns abstracted from just 500 prior examples. This training-free framework effectively bridges the gap between search-based and post-training methods, offering both performance enhancement and computational efficiency.
[…]
  • Broad Generalization Capability: Our framework demonstrates remarkable versatility across diverse tasks, domains, and modalities, validating the universality of high-level problem-solving thoughts across different problem spaces.
  • Plug-and-Play Flexibility: AStar’s modular design allows seamless integration with existing post-training strategies like GRPO, enhancing their capabilities without requiring structural modifications—particularly valuable in resource-constrained environments.
By introducing generalizable external reasoning guidelines that complement models’ internal capabilities, our approach provides a practical and scalable solution for enhancing multimodal reasoning capabilities even with limited computational resources. We anticipate that our findings will stimulate further exploration into the integration of explicit cognitive frameworks with MLLMs, advancing both theoretical understanding and practical applications across diverse domains.
Appendix B Preliminaries
This section describes Overall Notations (B.1), MLLM Reasoning (B.2), Monte Carlo Tree Search (B.3), and Verification Methods (Self-Consistency and Outcome Reward Model) (B.4).
B.1 Overall Notations
@@
B.2 MLLM Reasoning
With the advancement of computational resources and expanded datasets, MLLMs have demonstrated remarkable capabilities across various multimodal tasks [29, 1]. These range from visual reasoning tasks like chart understanding and visual question-answering [39] to more complex perception-action tasks in autonomous driving and robotics [68], where models must integrate visual inputs with decision-making processes. Central to these achievements is the development of effective reasoning methods, which can substantially enhance MLLM problem-solving capabilities, enabling even smaller models to achieve sophisticated reasoning abilities [45, 14, 27]. To formalize this reasoning process, we consider an autoregressive pre-trained MLLM serving as a policy model $\pi_{\theta}$. Given an input problem $x_{t}$, the model generates a sequence $output=(s_{0},s_{1},s_{2},...,s_{T})$ through iterative token prediction, where $s_{0}:=x_{t}$ represents the initial state and $s_{T}$ corresponds to the solution $y_{p}$. We define this generated sequence $(s_{0},s_{1},s_{2},...,s_{T})$ as a reasoning trajectory $y_{t}$. The conditional probability distribution of generating the complete reasoning trajectory is:
@@ B.2 MLLM Reasoning
Following previous studies, we can conceptualize MLLMs as world models, with the complex reasoning process formulated as a Markov decision process. Specifically, when addressing complex reasoning challenges in real-world scenarios, at each time step $t$, the model receives a state $S_{t-1}$, comprising the original input problem $x$ and preceding reasoning steps $(s_{0},s_{1},s_{2},...,s_{t-1})$. The policy model $\pi_{\theta}$ then generates the current action $a_{t}=\pi_{\theta}(\Phi(S_{t-1}))$, which prompts the MLLM to produce the next reasoning step $s_{t}$. The entire process, from the initial step $s_{0}$ to the final output $s_{T}$, naturally forms a complete trajectory or chain of thought.
@@
Table 6: Comparison with two primary reasoning methods: search-based methods (AR-MCTS, Mulberry) and post-training methods (SFT: LLaVA-CoT, LlamaV-o1, URSA; GRPO: LMM-R1, MM-Eureka, R1-VL). ‘OS Only’ indicates datasets constructed only with open-source models. AStar only requires 500 prior samples.
Method Type Open-Source Only Training-Free Training Data Training Cost $\downarrow$ Inference Cost $\downarrow$
AR-MCTS [14] Search 34.5K - High
Mulberry [15] Search 260K High Low
LLaVA-CoT [10] SFT 100K High Low
LlamaV-o1 [27] SFT 118K High Low
URSA [13] SFT 1100K High Low
LMM-R1-3B [20] GRPO 55.3K High Low
MM-Eureka-7B [48] GRPO 15K High Low
R1-VL-7B [18] GRPO 260K High Low
AStar (Ours) - 0.5K Low Low
@@
Early work explored multimodal chain-of-thought (CoT) [4], addressing the limitations of conventional MLLMs that typically defaulted to a simple “direct prediction” mode due to the scarcity of high-quality long-chain reasoning data. However, these approaches often exhibit instability in multimodal reasoning tasks due to the distribution shift between training and inference. Two primary methods have emerged to address this challenge: explicit search mechanisms and post-training methods. Recent advances are summarized in Table 6.
@@
(1) Explicit search mechanisms.   These approaches leverage explicit search structures (e.g., Monte Carlo tree search, MCTS) with specialized reward models to guide the exploration of solution paths. AR-MCTS [14] combines MCTS with active retrieval to expand the solution space and enhance performance, though its extensive iterations demand significant computational resources. Moreover, the quality and relevance of retrieved examples at each exploration step substantially impact model performance. Similarly, Mulberry [15] employs collective knowledge to enable collaborative conjecture, search, and identification of effective reasoning paths via MCTS. However, it requires substantial training data (260K samples) generated with expensive proprietary models, making it computationally intensive.
(2) Post-Training Methods.   This approach focus on developing structured reasoning capabilities through techniques such as Supervised Fine-Tuning (SFT) or Reinforcement Learning with Verifiable Rewards (RLVR). SFT-based approaches like LLaVA-CoT [10] and URSA [13] distill structured reasoning patterns through long-form CoT data, typically requiring substantial training data ($\geq$100K samples) and supervision from proprietary models. RLVR-based methods like LMM-R1 [20] and R1-VL [18] leverage techniques such as Group Relative Policy Optimization (GRPO) [17, 18], but encounter challenges with exploration depth and training stability. As shown in [22, 23], these methods primarily function by biasing the model’s output distribution toward reward-maximizing paths without introducing external knowledge, inherently limiting their exploration capacity.
@@
Figure 10: An illustration of the four phases in an iteration of MCTS for complex reasoning tasks.
B.3 Monte Carlo Tree Search
@@
Leveraging Monte Carlo Tree Search (MCTS), recent approaches like AR-MCTS [14] exploit MLLMs’ intrinsic capabilities and active retrieval for iterative exploration to enhance visual complex reasoning. However, these methods demand substantial computational resources due to extensive iterations. While approaches like Mulberry [15] attempt to integrate tree structures through model training, they require large-scale datasets and significant computational overhead. Moreover, current methods typically employ static reasoning processes, lacking the ability to adapt reasoning strategies based on problem complexity. In contrast, our approach employs MCTS only during the generation of prior reasoning patterns (referred to as "thought cards" in Sec. 2.2) and references these thought cards during inference to achieve efficient reasoning. This design enables AStar to adaptively match reasoning strategies to problem attributes, significantly reducing time complexity compared to traditional explicit search methods while maintaining comprehensive search coverage and performance. Furthermore, compared to traditional MCTS, we implement two key improvements in our card generation process to achieve efficient MCTS operations: (1) We introduce early termination strategies and integrate the MLLM’s reasoning process during MCTS iteration; (2) We use the likelihood of self-consistency majority voting as the reward value, enabling supervision-free generalization.
B.4 Verification Methods (Self-Consistency and Outcome Reward Model)
@@
After obtaining multiple valid candidate solutions for the test multimodal query $x_{t}=[q;i]$, selecting the most accurate reasoning trajectory among candidates presents a critical challenge. In multimodal reasoning, self-consistency represents a simple yet effective approach, where the final answer is determined through majority voting among sampled paths:
@@
Recent advances in LLM reasoning verification have demonstrated significant improvements in reasoning capabilities through verification modules [67]. This has motivated the integration of outcome and process supervision in multimodal domains [14, 13] to enhance reasoning performance. However, the scarcity of multimodal supervision data and challenges in ensuring supervision signal quality often compromise verification effectiveness, with few high-performance open-source verification models currently available. Therefore, for simplicity, we leverage mature text-domain outcome reward models (ORM).
Specifically, in the text domain, given a text problem $q$ and its reasoning trajectory $y_{t}$, ORM ($\mathcal{D}\times\mathcal{S}\rightarrow\mathbb{R}$) assigns a single real-value to indicate whether $y_{t}$ is correct. ORM is usually trained with a cross-entropy loss:
@@
where $y_{g}$ is the golden answer ($y_{g}=1$ if $traj$ is correct else $y_{g}=0$), and $r_{traj}$ is the sigmoid score of $traj$ assigned by ORM. The effectiveness of ORM heavily depends on training data quality. For reasoning problems with definitive answers, researchers can construct the ORM training set automatically by: (1) sampling candidate solutions, and (2) labeling solutions based on answer correctness. While this approach may introduce false positives (incorrect reasoning leading to correct answers), prior work has demonstrated its effectiveness in training robust ORM models [69]. This enables the development of high-performing open-source verification models in the text domain.
To leverage these powerful text-domain verification models for visual reasoning while maintaining our principle of balancing performance and efficiency, we first use an MLLM (e.g., Qwen2-VL-7B) to convert the visual input into a textual description, which is then combined with the original text query to form a pure-text question:
@@
where $\text{prompt}_{translate}$ denotes a translation prompt (e.g., “please translate the question and visual image into a pure text format”). This, along with the reasoning steps derived from AStar, is passed to the above ORM for evaluation. The top-3 highest-scoring paths are subsequently verified via self-consistency.
Note that, this process does not leak any heuristics towards specific tasks. The verification mechanism operates independently from the problem-solving process, maintaining task-agnosticity. The ORM evaluates solution quality based on general reasoning principles rather than task-specific shortcuts. Our experiments indicate that this approach achieves improvements over simple self-consistency with minimal computational overhead. Future work could explore integration with high-quality verification models for potentially better results. In this paper, we utilize off-the-shelf ORM model: Llama3.1-8B-ORM-Mistral-Data [^1] https://huggingface.co/RLHFlow/Llama3.1-8B-ORM-Mistral-Data.
Table 7: Comparison with other search-based LLM and MLLM methods. Note that, most methods contain limited space. In contrast, we define a rich set of reasoning actions, thus enhancing the upper bound of model reasoning capabilities.
@@
LLM Tree-of-Thought [70] $a_{3}$: one-step thought
RAP [71] $a_{5}$: divide and conquer
MLLM AR-MCTS [14] $a_{3}$: one-step thought
LLaVA-CoT [10] $a_{4}$: chain-of-thought
URSA [13] $a_{4}$: chain-of-thought
@@
Appendix C Algorithm Details
C.1 Action Space
Reasoning capabilities are fundamental for handling diverse tasks, encompassing multiple cognitive dimensions. Prior work has categorized reasoning into deductive, inductive, abductive, and analogical approaches [1]. While open-source MLLMs demonstrate basic task competency, advanced models necessitate more sophisticated human-like reasoning abilities to achieve their full potential. A key insight from recent studies suggests that model reasoning performance is bounded by their available reasoning actions [72]. Current frameworks, however, operate within a constrained action space (Table 7), which may limit the reasoning capabilities of MLLMs. To expand models’ reasoning capabilities beyond these constraints, we propose a comprehensive framework with six fundamental reasoning actions:
  • Visual Parsing ($a_{1}$): Specifically designed for visual reasoning, aimed at extracting and analyzing visual information to enable the model to integrate visual cues with textual reasoning
  • System Analysis ($a_{2}$): Incorporates complete image information and combines textual queries with image data to analyze task conditions and constraints
  • One-Step Thought ($a_{3}$): Generates the next one-step thought based on the given question (text and image information) and preceding reasoning steps.